A lot consists of 10 articles, 4 with minor defects, and 2 with major defects. One
article is chosen at random. Find the probability in the following cases.
i) "The item has no defects.
ii) The item has no major defects.
1
Expert's answer
2020-08-24T19:03:03-0400

Given that,

Total number of articles n = 10


No of articles with minor defects = 4


Probability of articles with minor defects Px = number of articles with minor defect ÷ total number of articles


Px = 4÷10 = 0.4


No of articles with major defects = 2


Probability of articles with minor defects Py = number of articles with minor defect ÷ total number of articles


Py = 2÷10 = 0.2


Answer 1)


On random selection of 1 article probability that there is no defect means there in no minor defect or no major defect is give by


P = 1 - (Px + Py) = 1 - Px - Py

P = 1 - 0.4 - 0.2 = 0.4

P = 0.4


Answer 2)


On random selection of 1 article probability that there is no Major defect (it can have a minor defect or completely good) is give by


P = 1 - Py = 1 - 0.2

P = 0.8